If A=(6i-8j) units, B=(-8i-3j) units, and C=(26i-19j) units, determine a and b such that aA+bB+C=0
1
Expert's answer
2020-05-20T09:32:10-0400

Given:

"A=6i-8j;\\
B=-8i-3j;\\
C=26i-19j."


"aA+bB+C=a(6i-8j)+b(-8i-3j)+26i-19j=0"

"\begin{matrix}
 6a-8b & =&-26 \\
 -8a-3b & =& 19
\end{matrix}"

Roots


"a=-\frac{115}{41},\quad b=\frac{47}{41}"